Susan can pick 4 pounds of coffee in an hour or gather 2 pounds of nuts. Tom can pick 2 pounds of coffee in an hour or gather 4 pounds of nuts. Each works 6 hours per day. Assume that Tom and Susan specialize according to their comparative advantage to produce coffee and nuts a. Suppose that Susan and Tom could buy or sell coffee and nuts in the world market at a price of $2 per pound for coffee and $2 per pound for nuts. If each person specialized completely in the good for which he or she had a comparative advantage, how much could they earn by selling all they produce (Toms earnings plus Susans earnings)? per day b. At the prices just described, what is the maximum amount of coffee Susan and Tom could buy in the world market with the income they eamed? The maximum amount of nuts? Maximum amount of coffee that could be bought: unds. Maximum amount of nuts that could be bought: punds. Using the world market and the prices des bed above, would it be possible for Susan and Tom to consume 40 pounds of nuts and 8 pounds of coffee each day, Click to select c. In light of their ability to buy and sell in world markets at the stated prices, on the same graph as their PPC, show all of the combinations of the two goods it would be possible for them to consume. This is also known as their consumption possibilities curve or CPC. Instruction: Use the tool provided to draw all combinations of coffee and nuts it would be possible for them to consume. Plot only the endpoints of this curveSusan & Toms PPC CPC 56 52 48 40 38 32 24 20 16 12 0 4 8 12 16 20 24 28 32 36 40 44 48 52 56 Nuts (pounds/day)
Answer #1

In the given question it has been said that Susan can pick 4 pounds of coffee or gather 2 pounds of nuts and Tom can pick 2 pounds of coffee or gather 4 pounds of nuts in every hour they will work. Now if they work for 6 hours in a day then Susan can pick total 24 pounds of coffe gather total 12 pounds of nuts. Similarly Tom can pick in total 6 hours 12 pounds of coffee or gather total 24 pounds of nuts.

Now Susan is specialised on picking coffee because in an hour Susan has to sacrifice less amount of nuts to pick one more coffee. Tom is specialised on gathering nuts because Tom has to sacrifice less amount of coffee to gather extra nut. So if Susan specialised in Coffee then Susan will pick total 24 pounds coffee in 6 hours and Tom will gather 24 pounfp nuts in 6 hours.

So in total in 6 hours they will able to pick 24 pounds coffee and will be able to gather 24 pounds of nuts.

a. If Susan and Tom could buy and sell coffee and nuts $2 per pound for coffee and $2 per pound for nut and if they produce according to their comparative advantage then Susan can earn $24*2=$48 and Tom can earn $2*24 =$48. So in total they will earn 48+48=$96 per day.

b. In the world market at the given price they can buy - Maximum amount of coffee that could be bought 96/2 (total earning/price) =48 pounds. Maximum amount of nuts that could be bought 96/2 =48 pounds.

In the given market price it will be possible to consume for Susan and Tom 40 pounds of nuts and 8 pounds of coffee. The answer is yes. Because in total they can spend $96. 40 pounds nuts price $80 and 8 pound coffee price $16. In total it becomes$96.

c. The counsumptco possibility curve and production possibility curve has been shown in following figure-

hD 24 28 24 12- PPe CPc